OpenStudy (anonymous):

What was the day of the week on 28th May, 2006?

OpenStudy (anonymous):

Please help out ...

Parth (parthkohli):

It was a Sunday (look at the calendar)

OpenStudy (anonymous):

Sunday :)

OpenStudy (anonymous):

sunday xD

OpenStudy (anonymous):

Yeah, sunday (look at the calendar too) :)

OpenStudy (anonymous):

how did you get the answer

Parth (parthkohli):

Wolframalpha xD

OpenStudy (anonymous):

I won't set a calender in the exam to look out

OpenStudy (anonymous):

I just went back through my calendar :) lol

OpenStudy (anonymous):

sunday

OpenStudy (anonymous):

Is there any mathematical logic that can be applied

Parth (parthkohli):

http://www.terra.es/personal2/grimmer/

Parth (parthkohli):

Here it is... the link

OpenStudy (anonymous):

hmm I don't know how to just figure that out then sorry :)

OpenStudy (anonymous):

oh great thanx....

OpenStudy (anonymous):

There is actually a way to do this.

OpenStudy (anonymous):

Mathematically I mean.

OpenStudy (anonymous):

pls.... pls provide an alternative way...also

OpenStudy (anonymous):

great @Directrix but I won't get a calender in the exam.

Parth (parthkohli):

http://www.terra.es/personal2/grimmer/ Again, see it..

Directrix (directrix):

Explanation: 28 May, 2006 = (2005 years + Period from 1.1.2006 to 28.5.2006) Odd days in 1600 years = 0 Odd days in 400 years = 0 5 years = (4 ordinary years + 1 leap year) = (4 x 1 + 1 x 2) 6 odd days Jan. Feb. March April May (31 + 28 + 31 + 30 + 28 ) = 148 days 148 days = (21 weeks + 1 day) 1 odd day. Total number of odd days = (0 + 0 + 6 + 1) = 7 0 odd day. Given day is Sunday. http://www.indiabix.com/aptitude/calendar/

OpenStudy (anonymous):

There are actually two way that I am aware of two years back I implemented this algorithm in C : http://en.wikipedia.org/wiki/Zeller%27s_congruence

OpenStudy (anonymous):

But this is not the way I would do now in a test ;)

OpenStudy (anonymous):

sunday

OpenStudy (anonymous):

@ParthKohli thanx for the link

OpenStudy (anonymous):

THANX GUYS....

OpenStudy (anonymous):

study more questions - http://www.job4india.in/questions-and-answers